# The CLAS Collaboration

### v2

@@ V.1 Filling kinematic cells with zero acceptance
Due to blind areas in the geometrical coverage of the CLAS detector, some kinematic bins of the double-pion production phase space turn out to have zero acceptance. In such bins, which are usually called empty cells, the cross section cannot be experimentally defined. These cells contribute to the integrals in Eqs. (14) along with all other kinematic bins. If ignored, the contribution from empty cells causes a systematic underestimation of the cross section and, therefore, some assumptions for their contents are needed. This situation causes some model dependence of the extracted cross sections.
The map of the empty cells was determined using the Monte Carlo simulation. A multi-dimensional cell was treated as empty if it contained generated events ($N_{\text{gen}}>$ 0), but did not contain any reconstructed events ($N_{\text{rec}}$ = 0). The cells with unreliable efficiencies, ruled out based on the 30% cut on the efficiency uncertainty, were also treated as empty (see Sec. IV.6).
@@
In the present work, empty multi-dimensional cells were filled with the Monte Carlo events generated with TWOPEG-D Skorodumina et al. (2017a). These events were subject to integral scaling, in order to adjust them to the experimental yield in the regular (non-empty) cells. The scaling was performed individually in each $\Delta W\Delta Q^{2}$ bin according to the ratio of the integrated yields of the experimental and reconstructed Monte Carlo events in the non-empty cells of that bin (see Refs. Skorodumina et al. (2020); Skorodumina for details).
Figure 9 introduces the single-differential cross sections given by Eqs. (14) and (15). The empty squares correspond to the case when the contribution from the empty cells was ignored, and the black circles are for the case when that was taken into account as described above. The figure indicates a satisfactory small contribution from the empty cells for the majority of data points (and therefore a small model dependence of the results). Only the edge points in the $\theta$ distributions (middle row) reveal pronounced contributions due to the negligible/zero CLAS acceptance in the corresponding directions.
For most of the $(W,Q^{2})$ points, the contribution from the empty cells to the fully integrated cross section is kept to a low level of $\sim$15%, slightly rising at the low $Q^{2}$ and high $W$ boundaries. Besides this, the empty cell contribution grows towards the threshold (i.e. for $W\lesssim 1.4$ GeV). This behavior was also observed in Refs. Fedotov et al. (2017, 2018, 2007, 2009) devoted to double-pion electroproduction off free protons.
To account for the model dependence, the approach established for the previous studies of double-pion cross sections was followed Isupov et al. (2017); Fedotov et al. (2017); Golovatch et al. (2019), i.e. the part of the single-differential cross section that came from the empty cells was assigned a 50% relative uncertainty (more details are in Sec. VI.2).
@@ V.2 Radiative corrections
The common way of handling this problem is to apply radiative corrections to the extracted cross sections. In this study, radiative corrections were performed using TWOPEG-D Skorodumina et al. (2017a), which is the event generator for the double-pion electroproduction off a moving proton. TWOPEG-D accounts for the radiative effects by means of the well-known Mo and Tsai approach Mo and Tsai (1969), which has traditionally been used for radiative corrections in studies of double-pion electroproduction Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009, 2017, 2018); Isupov et al. (2017); Trivedi and Gothe (2018). In Ref. Mo and Tsai (1969), the approach was applied to the inclusive case, while in TWOPEG-D, the fully integrated double-pion cross sections are used instead Skorodumina et al. (2017b, a).
In the employed approach Mo and Tsai (1969); Skorodumina et al. (2017b, a), the radiative photons are considered to be emitted collinearly either in the direction of the incoming or scattered electron (termed the “peaking approximation”). The calculation of the radiative cross section is split into two parts. The “soft” part assumes the energy of the emitted radiative photon to be less than a certain minimal value (10 MeV), while the “hard” part is for the photons with an energy greater than that value. The “soft” part is evaluated explicitly, while for the calculation of the “hard” part, an inclusive hadronic tensor is assumed. Based on previous experience, the latter assumption is considered to be adequate Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009, 2017, 2018); Isupov et al. (2017); Trivedi and Gothe (2018). Also note that approaches that are capable of describing radiative processes in exclusive double-pion electroproduction are not yet available.
Figure 10: Reciprocal of the radiative correction factor (i.e. $1/R$) as a function of $W$ for different $Q^{2}$ bins (see Eq. (17)).
The radiative correction factor $R$ in Eq. (10) was determined in the following way. Double-pion events either with or without radiative effects were generated with TWOPEG-D. Both radiated and non-radiated events were subjected to the Fermi smearing. Then the ratio given by Eq. (17) was taken in each $\Delta W\Delta Q^{2}$ bin.
@@
This approach gives the correction factor $R$ only as a function of $W$ and $Q^{2}$, disregarding its dependence on the hadronic variables. However, the need to integrate the cross section at least over four hadronic variables (see Eq. (14)) considerably reduces the influence of the final-state hadron kinematics on the radiative correction factor, thus justifying the applicability of the procedure Mo and Tsai (1969); Skorodumina et al. (2017b, a).
@@
V.3 Unfolding the effects of initial proton motion
In this study, information on the initial proton momentum is inaccessible for the majority of analyzed experimental events, as discussed above. For this reason, the invariant mass $W_{\textrm{i}}$ calculated under the target-at-rest assumption was used for the cross section binning (see Sec. IV.1), which leads to the Fermi smearing of the fully integrated and single-differential cross sections. The same reason necessitates the use of an approximate procedure of the Lab to CMS transformation (see Sec. IV.2). This approximation introduces some inaccuracy to the measured angular ($\theta$, $\varphi$, and $\alpha$) distributions, but does not affect the invariant mass distributions and $W$ and $Q^{2}$ cross section dependences due to the Lorentz invariance of these variables.
Being folded with the aforementioned effects of the initial proton motion, the extracted cross sections require a corresponding unfolding correction. This correction was performed by means of two Monte Carlo event generators TWOPEG Skorodumina et al. (2017b) and TWOPEG-D Skorodumina et al. (2017a). TWOPEG is the event generator for the double-pion electroproduction off a free proton that currently provides the best cross section estimation in the investigated kinematic region. TWOPEG-D is the event generator for the same exclusive reaction but off the proton that moves in the deuterium nucleus. This event generator was specifically developed for the studies where experimental information on the initial proton momentum is inaccessible and the target-at-rest assumption has to be employed. TWOPEG-D convolutes the double-pion cross section with effects of the initial proton motion, thus imitating the conditions of the experimental cross section extraction.
To calculate the correction factor, two samples of double-pion events, produced off protons at rest and off moving protons, were generated with TWOPEG and TWOPEG-D, respectively. For the latter, the smeared value of $W$ was used for the binning and the approximate Lab to CMS transformation was applied, so that the sample incorporated the same inaccuracies as the experimentally extracted cross sections.
@@ V.3 Unfolding the effects of initial proton motion
where $\mathbb{N}_{\text{nofermi}}$ and $\mathbb{N}_{\text{fermi}}$ are the weighted numbers of generated double-pion events in the $\Delta W\Delta Q^{2}\Delta^{5}\tau$ bin produced off the proton at rest and off the moving proton, respectively.
For most points of the fully integrated cross sections at $W>$ 1.4 GeV, the correction does not exceed ±$\pm 5$%. However, for $W$ closer to the threshold, the correction grows rapidly, reaching up to about 40% for the first $W$ bin (see Refs. Skorodumina et al. (2020); Skorodumina for details).
The value of the correction factor in Eq. (18) depends on both the free proton cross sections and the model of the deuteron wave function implemented into the event generators. The former relies strongly on the JM model fit of the available data on the double-pion cross sections, while for the latter, the Bonn model was used (see Refs. Skorodumina et al. (2017b, a)). Therefore, the uncertainty of the extracted cross sections that comes from the unfolding correction was attributed to the model-dependent uncertainty as discussed in Sec. VI.2.
@@ V.4 Corrections for binning effects
In general, being extracted in a finite bin, the cross section naturally undergoes averaging within this bin, and this averaged value is then assigned to the bin central point. Any nonlinear behavior of the cross section within the bin will likely result in an offset of the obtained cross section value. To cure this effect, a binning correction was applied that includes a cubic spline approximation for the cross section shape Skorodumina et al. (2020); Skorodumina . Due to the relatively fine binning in all kinematic variables used in this study, the influence of the binning effects on the cross sections is marginal: the typical value of the correction is $\sim$1rising up to 5% for some data points at low $W$.
In addition to that, the cross section in the next to last point of the invariant mass distributions was subject to a specific separate correction, which is described in detail in Refs. Fedotov et al. (2017, 2018); Skorodumina et al. (2020); Skorodumina . The need for this correction follows from the broadening of the reaction phase space with $W$, which causes the upper boundary of the invariant mass distributions to be $W$ dependent.
@@ VI Cross section uncertainties
In this study (like in other studies of the double-pion cross sections Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018)) three separate types of cross section uncertainties were considered, i.e. statistical uncertainties, uncertainties due to the model dependence, and systematic uncertainties.
@@ VI.1 Statistical uncertainties
The absolute statistical uncertainty due to the limited statistics of the experimental data was calculated in the non-empty multi-dimensional bins as
$\delta_{\text{stat}}^{\text{exp}}(\Delta^{5}\tau)=\frac{1}{\mathcal{E}\!\cdot\!R\!\cdot\!\mathcal{F}\!\cdot\!\Gamma_{\text{v}}}\cdot\frac{\sqrt{\left(\frac{N_{\text{full}}}{Q_{\text{full}}^{2}}+\frac{N_{\text{empty}}}{Q_{\text{empty}}^{2}}\right)}}{\Delta W\!\cdot\!\Delta Q^{2}\!\cdot\!\Delta^{5}\tau\!\cdot\!\left[\mathcal{L}\right]},$ (19)
@@
The absolute uncertainty due to the limited Monte Carlo statistics was estimated in the non-empty bins as
@@
where $\frac{\textrm{d}^{5}\sigma_{\text{v}}}{\textrm{d}^{5}\tau}$ is the virtual photoproduction cross section given by Eq. (11), $\mathcal{E}$ is the efficiency inside the multi-dimensional bin defined by Eq. (16), and $\delta\mathcal{E}$ is the absolute statistical efficiency uncertainty.
The calculation of the efficiency uncertainty $\delta\mathcal{E}$ is not straightforward because (i) $N_{\text{gen}}$ and $N_{\text{rec}}$ in Eq. (16) are not independent and (ii) Monte Carlo events in this equation are subject to weighting. Therefore, a special approach described in Ref. Laforge and Schoeffel (1997) was used to calculate $\delta\mathcal{E}$. Neglecting the event migration between bins, this approach gives the following expression for the absolute statistical uncertainty of the efficiency in a bin for the case of a weighted Monte Carlo simulation,
@@
where $N_{\text{gen}}$ and $N_{\text{rec}}$ are the number of the generated and reconstructed Monte Carlo events inside the multi-dimensional bin, respectively, $\mathbb{N}_{\text{gen}}$ and $\mathbb{N}_{\text{rec}}$ are the corresponding weighted event numbers, and $w$ is the weight of an individual event.
The two parts of the statistical uncertainty given by Eqs. (19) and (20) were combined quadratically into the total absolute statistical uncertainty in each non-empty $\Delta^{5}\tau$ bin,
$\delta_{\text{stat}}^{\text{tot}}(\Delta^{5}\tau)=\sqrt{\left(\delta_{\text{stat}}^{\text{exp}}\right)^{2}+\left(\delta_{\text{stat}}^{\text{MC}}\right)^{2}}.$ (22)
@@
For the extracted single-differential cross sections, the statistical uncertainty $\delta_{\text{stat}}^{\text{tot}}(\Delta X)$ (where $X$ is one of the final state variables, e.g. $M_{h_{1}h_{2}}$, $\theta_{h_{1}}$, $\alpha_{h_{1}}$) was obtained from the uncertainties $\delta_{\text{stat}}^{\text{tot}}(\Delta^{5}\tau)$ of the five-fold differential cross sections according to the standard error propagation rules.
@@ VI.2 Model-dependent uncertainties
In studies of double-pion production off free protons, the cross section model dependence originates from the filling of the empty cells and the corresponding cross section uncertainty is commonly treated as a separate uncertainty type Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018). In this analysis, one further source of the model dependence appears, which is the correction that unfolds the effects of the initial proton motion (see Sec. V.3). The two sources were found to give comparable uncertainties for the two lowest $W$ bins, while for the other bins the dominant part of the model-dependent uncertainty comes from the filling of the empty cells.
Both the contribution from the empty cells and the value of the unfolding correction vary greatly (from insignificant to considerable) for different bins in the final hadronic variables. Therefore, the model-dependent uncertainties were estimated in each $\Delta X$ bin of the single-differential cross sections (where $X$ is one of the final state variables introduced in Sec. IV.3).
The absolute cross section uncertainty $\delta^{\text{cells}}_{\text{model}}(\Delta X)$ due to the filling of the empty cells was calculated by
@@
where the parentheses contain the difference between the cross section values calculated with the empty cell contributions (“filled”) and without them (“not filled”), see also Fig. 9.
For each $\Delta X$ bin of the single-differential distributions, the relative uncertainty due to the unfolding procedure was estimated by

### v4

@@ V.1 Filling kinematic cells with zero acceptance
Due to blind areas in the geometrical coverage of the CLAS detector, some kinematic bins of the double-pion production phase space turned out to have zero acceptance. In such bins, which are usually called empty cells, the cross section cannot be experimentally defined. These cells contribute to the integrals in Eqs. (14) along with all other kinematic bins. If ignored, the contribution from the empty cells causes a systematic underestimation of the cross section, and therefore some assumptions for their contents are needed. This situation causes some model dependence of the extracted cross sections.
The map of the empty cells was determined using the Monte Carlo simulation. A multi-dimensional cell was treated as empty if it contained generated events ($N_{\text{gen}}>$ 0) but did not contain any reconstructed events ($N_{\text{rec}}$ = 0). The cells with unreliable efficiencies, ruled out based on the 30% cut on the efficiency uncertainty, were also treated as empty (see Sec. IV.6).
@@
In the present work, empty multi-dimensional cells were filled with generated Monte Carlo events produced by TWOPEG-D Skorodumina et al. (2017a). These events were subject to integral scaling, in order to adjust their yield to the yield of experimental events in regular (nonempty) cells. The scaling was performed individually in each $\Delta W\Delta Q^{2}$ bin according to the ratio of the integrated yields of experimental events and reconstructed Monte Carlo events in all nonempty cells of that bin (see Refs. Skorodumina et al. (2020); Skorodumina for details).
Figure 9 introduces the single-differential cross sections given by Eqs. (14) and (15). The open squares correspond to the case when the contribution from the empty cells was ignored, and the black circles are for the case when it was taken into account as described above. The figure indicates a satisfactory small contribution from the empty cells for the majority of data points (and therefore a small model dependence of the results). Only the edge points in the $\theta$ distributions (middle row) reveal pronounced contributions due to the negligible or zero CLAS acceptance in the corresponding directions.
For most of the $(W,Q^{2})$ points, the contribution from the empty cells to the fully integrated cross section is kept to a low level of $\approx$15%, slightly rising at the low-$Q^{2}$ and high-$W$ boundaries. Besides this, the empty cells contribution grows towards the reaction threshold (i.e. for $W\lesssim 1.4$ GeV). This behavior was also observed in Refs. Fedotov et al. (2017, 2018, 2007, 2009) devoted to double-pion electroproduction off free protons.
To account for the model dependence, the approach established in previous studies of double-pion cross sections was followed Isupov et al. (2017); Fedotov et al. (2017); Golovatch et al. (2019), i.e. the part of the single-differential cross section that came from the empty cells was assigned a 50% relative uncertainty (more details are in Sec. VI.2).
@@ V.2 Radiative corrections
The common way of handling this problem is to apply radiative corrections to the extracted cross sections. In this study, radiative corrections were performed using TWOPEG-D Skorodumina et al. (2017a), which is an event generator for double-pion electroproduction off a moving proton. TWOPEG-D accounts for radiative effects by means of the well-known Mo and Tsai approach Mo and Tsai (1969), which has traditionally been used for radiative corrections in studies of double-pion cross sections Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009, 2017, 2018); Isupov et al. (2017); Trivedi and Gothe (2018). In Ref. Mo and Tsai (1969), the approach was applied to the inclusive case, while in TWOPEG-D, the fully integrated double-pion cross sections are used instead Skorodumina et al. (2017b, a).
In the employed approach Mo and Tsai (1969); Skorodumina et al. (2017b, a), the radiative photons are considered to be emitted collinearly either in the direction of the incoming or scattered electron (termed the “peaking approximation”). The calculation of the radiative cross section is split into two parts. The “soft” part assumes the energy of the emitted radiative photon to be less than a certain minimal value (10 MeV), while the “hard” part is for the photons with an energy greater than that value. The “soft” part is evaluated explicitly, while for the calculation of the “hard” part, an inclusive hadronic tensor is assumed. Based on previous experience with radiative corrections for double-pion cross sections, the latter assumption is considered to be adequate Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009, 2017, 2018); Isupov et al. (2017); Trivedi and Gothe (2018). Also note that approaches that are capable of describing radiative processes in exclusive double-pion electroproduction are not yet available.
Figure 10: Reciprocal of the radiative correction factor (i.e. $1/R$) as a function of $W$ for different $Q^{2}$ bins [see Eq. (17)].
The radiative correction factor $R$ in Eq. (10) was determined in the following way. Double-pion events either with or without radiative effects were generated with TWOPEG-D. Both radiated and nonradiated events were subjected to the Fermi smearing. Then the ratio given by Eq. (17) was taken in each $\Delta W\Delta Q^{2}$ bin.
@@
This approach gives the correction factor $R$ only as a function of $W$ and $Q^{2}$, disregarding its dependence on the hadronic variables. However, the need to integrate the cross section at least over four hadronic variables [see Eq. (14)] considerably reduces the influence of the final-state hadron kinematics on the radiative correction factor, thus justifying the applicability of the procedure Mo and Tsai (1969); Skorodumina et al. (2017b, a).
@@
V.3 Unfolding the effects of the initial proton motion
In this study, information on the initial proton momentum is inaccessible for the majority of analyzed experimental events, as discussed above. For this reason, the invariant mass $W_{\textrm{i}}$ calculated under the target-at-rest assumption was used for the cross section binning (see Sec. IV.1), which leads to the Fermi smearing of the fully integrated and single-differential cross sections. The same reason necessitates the use of an approximate procedure of the Lab-to-CMS transformation (see Sec. IV.2). This approximation introduces some inaccuracy to the measured angular ($\theta$, $\varphi$, and $\alpha$) distributions but does not affect the invariant mass distributions and $W$ and $Q^{2}$ cross section dependences due to the Lorentz invariance of these variables.
Being folded with the aforementioned effects of the initial proton motion, the extracted cross sections require a corresponding unfolding correction. This correction was performed by means of two Monte Carlo event generators TWOPEG Skorodumina et al. (2017b) and TWOPEG-D Skorodumina et al. (2017a). TWOPEG is an event generator for double-pion electroproduction off a free proton that currently provides the best cross section estimation in the investigated kinematic region. TWOPEG-D is an event generator for the same exclusive reaction but off a proton that moves in a deuterium nucleus. This event generator was specifically developed for the studies where experimental information on the initial proton momentum is inaccessible and the target-at-rest assumption has to be employed. TWOPEG-D convolutes double-pion cross sections with the effects of the initial proton motion, thus imitating the conditions of experimental cross section extraction.
To calculate the correction factor, two samples of double-pion events, produced off protons at rest and off moving protons, were generated with TWOPEG and TWOPEG-D, respectively. For the latter, the smeared value of $W$ was used for the binning and the approximate Lab-to-CMS transformation was applied, so that the sample incorporated the same inaccuracies as the experimentally extracted cross sections.
@@ V.3 Unfolding the effects of the initial proton motion
where $\mathbb{N}_{\text{noFermi}}$ and $\mathbb{N}_{\text{Fermi}}$ are the weighted numbers of generated double-pion events in a $\Delta W\Delta Q^{2}\Delta^{5}\tau$ bin produced off protons at rest and off moving protons, respectively.
For most points of the fully integrated cross sections at $W>$ 1.4 GeV, the correction does not exceed ±$\pm 5$%. However, for $W$ closer to the reaction threshold, the correction grows rapidly, reaching up to about 40% for the first $W$ bin (see Refs. Skorodumina et al. (2020); Skorodumina for details).
The value of the correction factor in Eq. (18) depends on both the free proton cross sections and the deuteron wave function implemented into the event generators. The former relies strongly on the JM model fit of the available data on double-pion cross sections, whereas for the latter, the Bonn model was used (see Refs. Skorodumina et al. (2017b, a)). Therefore, the uncertainty of the extracted cross sections that comes from the unfolding correction was attributed to the model-dependent uncertainty as discussed in Sec. VI.2.
@@ V.4 Corrections for binning effects
Being extracted in a finite bin, the cross section naturally undergoes averaging within this bin. In the present work, this averaged value is then assigned to the bin central point. Any nonlinear behavior of the cross section within the bin will likely result in an offset of the obtained cross section value. To cure this effect, binning corrections were applied to the fully integrated and single-differential cross sections that included a cubic spline approximation for the cross section shape Skorodumina et al. (2020); Skorodumina . Due to the relatively fine binning in all kinematic variables used in this study, the influence of binning effects on the cross sections is marginal: the typical value of the correction is $\approx$1rising up to 5% for some data points at low $W$.
In addition to that, the cross section in the next-to-last point of the invariant mass distributions was subject to a separate correction, which is described in detail in Refs. Fedotov et al. (2017, 2018); Skorodumina et al. (2020); Skorodumina . The need for this correction follows from the broadening of the reaction phase space with $W$, which causes the upper boundary of the invariant mass distributions to be $W$ dependent.
@@ VI Cross section uncertainties
In this study (like in other studies of double-pion cross sections Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018)) three separate types of cross section uncertainties were considered, i.e. statistical uncertainties, uncertainties due to the model dependence, and systematic uncertainties.
@@ VI.1 Statistical uncertainties
The absolute statistical uncertainty due to the limited statistics of the experimental data was calculated in a nonempty multi-dimensional bin as
$\delta_{\text{stat}}^{\text{expt}}(\Delta^{5}\tau)=\frac{1}{\mathcal{E}\!\cdot\!R\!\cdot\!\mathcal{F}\!\cdot\!\Gamma_{\text{v}}}\cdot\frac{\sqrt{\left(\frac{N_{\text{full}}}{Q_{\text{full}}^{2}}+\frac{N_{\text{empty}}}{Q_{\text{empty}}^{2}}\right)}}{\Delta W\!\cdot\!\Delta Q^{2}\!\cdot\!\Delta^{5}\tau\!\cdot\!\left[\mathcal{L}\right]},$ (19)
@@
The absolute uncertainty due to the limited Monte Carlo statistics was estimated in nonempty bins as
@@
where $\frac{\textrm{d}^{5}\sigma_{\text{v}}}{\textrm{d}^{5}\tau}$ is the virtual photoproduction cross section given by Eq. (11), $\mathcal{E}$ is the efficiency inside a multi-dimensional bin defined by Eq. (16), and $\delta\mathcal{E}$ is the absolute efficiency uncertainty.
The calculation of the efficiency uncertainty $\delta\mathcal{E}$ is not straightforward because (i) $N_{\text{gen}}$ and $N_{\text{rec}}$ in Eq. (16) are not independent and (ii) Monte Carlo events in this equation are subject to weighting. Therefore, a special approach described in Ref. Laforge and Schoeffel (1997) was used to calculate $\delta\mathcal{E}$. Neglecting event migration between bins, this approach gives the following expression for the absolute statistical uncertainty of the efficiency in a bin for the case of a weighted Monte Carlo simulation:
@@
where $N_{\text{gen}}$ and $N_{\text{rec}}$ are the numbers of generated and reconstructed Monte Carlo events inside a multi-dimensional bin, respectively, $\mathbb{N}_{\text{gen}}$ and $\mathbb{N}_{\text{rec}}$ are the corresponding weighted event numbers, and $w$ is the weight of an individual event.
The two parts of the statistical uncertainty given by Eqs. (19) and (20) were combined quadratically into the total absolute statistical uncertainty in each nonempty $\Delta^{5}\tau$ bin,
$\delta_{\text{stat}}^{\text{tot}}(\Delta^{5}\tau)=\sqrt{\left(\delta_{\text{stat}}^{\text{expt}}\right)^{2}+\left(\delta_{\text{stat}}^{\text{MC}}\right)^{2}}.$ (22)
@@
For the extracted single-differential cross sections, the statistical uncertainty $\delta_{\text{stat}}^{\text{tot}}(\Delta X)$ was obtained from the uncertainty $\delta_{\text{stat}}^{\text{tot}}(\Delta^{5}\tau)$ of the five-fold differential cross section according to the standard error propagation rules (where $X$ denotes one of the final-state variables, e.g. $M_{h_{1}h_{2}}$, $\theta_{h_{1}}$, $\alpha_{h_{1}}$).
@@ VI.2 Model-dependent uncertainties
In studies of double-pion production off free protons, the cross section model dependence originates from the filling of the empty cells, and the corresponding cross section uncertainty is commonly treated as a separate uncertainty type Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018). In this analysis, one further source of the model dependence appears, which is the correction that unfolds the effects of the initial proton motion (see Sec. V.3). The two sources were found to give comparable uncertainties for the two lowest $W$ bins, while for the other bins the dominant part of the model-dependent uncertainty comes from the empty cells filling.
Both the contribution from the empty cells and the value of the unfolding correction vary greatly (from insignificant to considerable) for different bins in the final hadronic variables. Therefore, the model-dependent uncertainties were estimated in each $\Delta X$ bin of the single-differential cross sections (where $X$ is one of the final-state variables introduced in Sec. IV.3).
The absolute uncertainty $\delta^{\text{cells}}_{\text{model}}(\Delta X)$ due to the filling of the empty cells was calculated by
@@
where the parentheses contain the difference between the cross section values calculated with the empty cells contributions (“filled”) and without them (“not filled”). See also Fig. 9.
For each $\Delta X$ bin of the single-differential distributions, the relative uncertainty due to the unfolding correction was estimated by